OpenStudy (anonymous):

simplify the radical expression 3 square root of 5+6 square root of 45

OpenStudy (anonymous):

what this would look like in math would be \[3\sqrt{5+6}\sqrt{45}\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so what you do is simplify the things under the radical sign first. That would make it now \[3\sqrt{11}\sqrt{45}\]

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 3\sqrt{11} ~~\sqrt{45} }\) \(\LARGE\color{blue}{ \bf 3\sqrt{11} ~3\sqrt{5} }\) \(\LARGE\color{blue}{ \bf 3(3)\sqrt{11(5)} }\) \(\LARGE\color{blue}{ \bf 9\sqrt{55} }\)

OpenStudy (anonymous):

What you would say here would be 3 times the square root of 11 times the square root of 45. It doesn't matter what you multiply first, as long as you multiply all of them. So after you compute all of it, the answer, rounded to the nearest tenth, is 66.7

OpenStudy (anonymous):

or that to :/ and his answer is right because it asks you to "simplify" not solve like i did.

OpenStudy (anonymous):

yet you can simplify further than that right solomon?

OpenStudy (anonymous):

\[9*11\sqrt{5}?\]

OpenStudy (anonymous):

then the answer is 99sqrt5

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No problem.